# REST API Testing in DevOps: A Study on an Evolving Healthcare IoT Application

### v1

@@ 4.6.2 Internal validity threat
The threat to interval validity can occur due to uncontrolled factors. A possible threat to the internal validity of our experiments may occur due to any instrumentation performed on the REST API testing tools or the REST APIs under test. To mitigate this threat, we selected REST API testing tools in their executable form, without modifying their code. In the case of REST APIs under test, we only have black-box access to these APIs, with all implementations carried out by the industry partner. Another possibility of internal validity threat can be due to tool inputs, including the API schema and test configuration, particularly when using RESTest. To reduce this threat, we relied on the documentation provided for each REST API testing tool to ensure proper setup and usage. Additionally, we conducted multiple sessions with the technical team from Oslo City and their industry partner. The primary purpose of these sessions was to demonstrate the setup process and acquire their feedback.
@@ 4.6.3 Construct validity threat
Construct validity threat is associated with the interpretation of results. We presented our evaluation findings using metrics commonly employed in similar experimental contexts [27, 9, 48]. Another potential threat to construct validity could arise from the manual analysis of faults and regressions. To mitigate this threat, we shared our experimental data, which includes generated test cases, failure reports, test coverage, regressions, and faults catalog, with Oslo City’s healthcare department for their review and feedback. Due to the sensitive nature of the data involved in this study, relating to API information and patient-associated data, we are bound by non-disclosure agreements that prevent us from publicly releasing a replication package or data set.
@@ 4.6.4 Conclusion validity threat
The threat to conclusion validity arises from the selection of analysis metrics, measures, or statistical tests that may influence the accurate drawing of conclusions. A potential threat to the validity of the conclusions of experiments may arise due to the randomness factor of REST API testing tools. We ran each tool 10 times with a fixed one-hour duration to address the randomness factor. This experimental setting is aligned with existing empirical studies [9, 48]. To analyze and report the results of our experiments, we employed metrics such as API failures, faults, test coverage, and regressions in evolving releases, which are commonly used in similar studies [27, 12]. Additionally, we evaluated the results using non-parametric statistical tests, specifically the Wilcoxon signed-rank test and Vargha-Delaney Â12 effect size measure, following recommended guidelines [60].
Figure 7: A workflow illustrating the recommendations for practitioners in selecting REST API testing tools.
@@ 5 Tool Selection Recommendations
Based on our experimental findings, we provide recommendations for practitioners to assist in selecting REST API testing tools, taking into account the testing goal and cost overhead. Figure 7 presents a workflow serving as a visual guide to illustrate the selection process. As observed in experimental results, there is no one-size-fits-all tool that is universally effective for all testing purposes. The performance and effectiveness of each tool can vary depending on the specific testing context. Therefore, testers need to define the testing goal, i.e., whether they are interested in analyzing failures, faults, coverage, regressions, or a combination of these aspects. In the following, we discuss recommendations considering the testing goal and cost overhead.
  • If the testing goal is to find the maximum number of unique failures and a high-cost overhead is acceptable, EvoMaster-BB can be a suitable choice. However, if the goal is to identify as many unique failures as possible with minimum cost overhead, RESTest-FT seems to be a reasonable alternative.
@@
  • If the testing goal is to identify the maximum number of unique failures and faults, and high-cost overhead is not a concern, one can opt for RESTest-CBT or RESTest-RT. However, if the requirement is to minimize cost overhead while detecting a higher number of unique failures and faults, RESTest-FT can be an appropriate choice.
  • When the testing goal is to identify the maximum number of unique failures while achieving high coverage, and if high-cost overhead is acceptable, EvoMaster-BB can be a suitable option. Conversely, if there is a need to reduce cost overhead, RestTestGen-NET can be viewed as a viable alternative.
@@
  • If testing aims to uncover a maximum number of unique failures and regressions without high-cost overhead constraints, EvoMaster-BB can be an appropriate selection. Conversely, if the intent is to reduce cost overhead while maintaining the same testing goal, Schemathesis-PT can be considered a suitable option.
@@ 6.1 Language-specific Test Inputs
The healthcare application used in our experiments operates in several countries and provides multilingual support. This trait is common among web services of multinational companies, which typically support multiple languages. An important aspect of testing REST APIs of such applications is using test inputs in all supported languages, necessitating the generation of language-specific test inputs. For instance, if an application supports Norwegian, Swedish, and English, generating test inputs in these languages is needed to test the application’s APIs from the language perspective. Moreover, if an application is region-specific with API supporting a particular language, it would require testing tools to generate language-specific inputs. For example, the cancer registry of Norway receives and compiles cancer cases in the Norwegian language [63]. In this case, the testing tool must generate test inputs in Norwegian.
@@
6.2 API Keys and Access Tokens
Healthcare IoT applications, such as the one in our study, are integrated with different third-party applications using API keys and access tokens for authentication and authorization. Each of these third-party applications supports different types of tokens (for example, OAuth2) with a certain level of security. Generating test cases for APIs that require access to third-party applications involves generating API keys and access tokens, which must adhere to a specific format. The automatic generation of valid API keys and access tokens presents a significant challenge.
@@ 6.3 Automation Level
Each REST API testing tool supports different automation levels. In addition to adhering to the OpenAPI specification, some tools require additional manual effort. For instance, RESTest requires writing test configurations. This effort escalates considerably as the REST APIs and their specifications evolve during the DevOps process. A potential solution to alleviate manual effort involves utilizing automated tools that can infer REST API specifications as their implementation undergoes evolution. For example, a tool like RESTSpecIT [64] can be used for this purpose.
@@ 6.4 Regression Test Selection, Minimization, and Prioritization
Numerous techniques are presented in the literature for prioritization, minimization, and selection of the test suite in the context of regression testing [53, 54]. However, to the best of our knowledge, their applicability to prioritize, minimize, and select regression test cases for REST APIs remains undetermined. Proposing new techniques for regression test suite prioritization, minimization, or selection falls outside the scope of this paper. Therefore, in our experiments, we used retest-all strategy, which involves running all the tests generated for the initial release on all subsequent releases. Given the daily or weekly release routine within the DevOps process, executing all test cases becomes impractical due to the significant overhead of API calls. For IoT applications’ REST APIs dependent on various third-party applications, the frequent execution of numerous tests is infeasible due to the API call limit imposed by these external applications. A potential area for future research could focus on devising strategies for prioritization, minimization, or selection of test cases specifically for the regression testing of REST APIs of IoT-based applications. The regression testing outcomes presented in this work could serve as a foundation for further research. These findings could also be used to refine or customize existing methods for test selection, minimization, and prioritization within the REST API testing context.
@@ 7 Conclusion
The results revealed that EvoMaster-BB generated the highest number of test cases compared to all other tools. On the other hand, RestTestGen-MAST did not generate any tests, indicating that none of the APIs has security vulnerabilities. In terms of API failures, all tools generated a higher number of client-side errors than server-side errors. Among all, EvoMaster-BB generated the highest number of errors on the client and server side. The tests produced by all tools led to the identification of 18 unique potential faults across 14 releases, with RESTest-CBT and RESTest-RT being the primary contributors to these faults. As for API test coverage, the results indicated that all tools attained a coverage level of 82-84%, with the lowest coverage observed in status code and status code class coverages. The regression results showed that tests generated by EvoMaster-BB led to finding the highest number of regressions. In total, 23 unique potential regressions were identified in all 14 releases. Finally, the cost analysis demonstrated that each tool generates a significant number of tests, often leading to numerous duplicate failures, thereby resulting in a cost overhead exceeding 80%.

### v2

@@ 4.6.2 Internal validity threat
The threat to interval validity can occur due to uncontrolled factors. A possible threat to the internal validity of our experiments may occur due to any instrumentation performed on the REST API testing tools or the REST APIs under test. To mitigate this threat, we selected REST API testing tools in their executable form, without modifying their code. In the case of REST APIs under test, we only have black-box access to these APIs, with all implementations carried out by the industry partner. Another possibility of internal validity threat can be due to tool inputs, including the API schema and test configuration, particularly when using RESTest. To reduce this threat, we relied on the documentation provided for each REST API testing tool to ensure proper setup and usage. Furthermore, we conducted multiple sessions with the technical team from Oslo City and their industry partner. The primary purpose of these sessions was to demonstrate the setup process and acquire their feedback. An additional potential threat could arise from the parallel execution of tools, which may dominate CPU usage and result in unequal resource distribution. To handle this, we adopted a parallel execution setup inspired by existing work of a similar nature zhang2023open . Following that setup, we configured multiple threads to execute parallel jobs while efficiently utilizing the available logical processors and CPU cores.
@@ 4.6.3 Construct validity threat
Construct validity threat is associated with the interpretation of results. We presented our evaluation findings using metrics commonly employed in similar experimental contexts martin2022online ; kim2022automated ; zhang2023open . Another potential threat to construct validity could arise from the manual analysis of faults and regressions. To mitigate this threat, we shared our experimental data, which includes generated test cases, failure reports, test coverage, regressions, and fault catalogs, with Oslo City’s healthcare department for their review and feedback. Due to the sensitive nature of the data involved in this study, relating to API information and patient-associated data, we are bound by non-disclosure agreements that prevent us from publicly releasing a replication package or data set.
@@ 4.6.4 Conclusion validity threat
The threat to conclusion validity arises from the selection of analysis metrics, measures, or statistical tests that may influence the accurate drawing of conclusions. A potential threat to the validity of the conclusions of experiments may arise due to the randomness factor of REST API testing tools. We ran each tool 10 times with a fixed one-hour duration to address the randomness factor. This experimental setting is aligned with existing empirical studies kim2022automated ; zhang2023open . To analyze and report the results of our experiments, we employed metrics such as API failures, faults, test coverage, and regressions in evolving releases, which are commonly used in similar studies martin2022online ; godefroid2020differential . Moreover, we evaluated the results using non-parametric statistical tests, specifically the Wilcoxon signed-rank test, Vargha-Delaney Â12 effect size measure, and Holm’s correction to adjust p-values, following recommended guidelines arcuri2011practical .
Figure 9: Workflow illustrating recommendations for practitioners in selecting REST API testing tools for healthcare applications.
@@ 5 Tool Selection Recommendations
Based on our experimental findings with testing the REST API of an industrial healthcare application, we provide recommendations to guide practitioners in selecting suitable REST API testing tools for similar healthcare systems. Figure 9 presents a workflow serving as a visual guide to illustrate the selection process. As observed in the experimental results, there is no one-size-fits-all tool that is universally effective for all testing purposes. The performance and effectiveness of each tool can vary depending on the specific testing context. Therefore, testers need to define the testing goal, i.e., whether they are interested in analyzing failures, faults, coverage, regressions, or a combination of these aspects. In the following, we discuss recommendations considering the testing goal and cost overhead.
  • If the testing goal is to find the maximum number of failures and a high-cost overhead is acceptable, EvoMaster-BB can be a suitable choice. However, if the goal is to identify as many failures as possible with minimum cost overhead, RESTest-FT seems to be a reasonable alternative.
@@
  • If the testing goal is to identify the maximum number of failures and faults, and high-cost overhead is not a concern, one can opt for RESTest-CBT or RESTest-RT. However, if the requirement is to minimize cost overhead while detecting a higher number of failures and faults, RESTest-FT can be an appropriate choice.
  • When the testing goal is to identify the maximum number of failures while achieving high coverage, and if high-cost overhead is acceptable, EvoMaster-BB can be a suitable option. Conversely, if there is a need to reduce cost overhead, RestTestGen-NET can be viewed as a viable alternative.
@@
  • If testing aims to uncover a maximum number of failures and regressions without high-cost overhead constraints, EvoMaster-BB can be an appropriate selection. Conversely, if the intent is to reduce cost overhead while maintaining the same testing goal, Schemathesis-PT can be considered a suitable option.
@@
All tool recommendations outlined above are formulated by carefully considering key testing-related aspects, such as cost and coverage. In the Oslo City context, we shared all results, findings, and recommendations with their team and industry partner, both during in-person meetings and through this paper. While they found these insightful, the decision to select a specific tool remained entirely theirs. Our role was to provide tool configurations, test results, and a selection guide to support their decision-making based on their testing needs and resource constraints. It is also important to note that our objective was not to conduct an industrial evaluation of testing tools with practitioners, but to analyze their effectiveness for the given REST API of the healthcare IoT application.
@@ 6.1 Language-specific Test Inputs
The healthcare application used in our experiments operates in several countries and provides multilingual support. This trait is common among web services of multinational companies, which typically support multiple languages. An important aspect of testing REST APIs of such applications is using test inputs in all supported languages, necessitating the generation of language-specific test inputs. For instance, if an application supports Norwegian, Swedish, and English, generating test inputs in these languages is needed to test the application’s APIs from a language perspective. Moreover, if an application is region-specific with an API supporting a particular language, it would require testing tools to generate language-specific inputs. For instance, in our experiments with the Karie and Medido device APIs, we found that when the language setting was Norwegian, the APIs expected inputs in Norwegian. When REST API testing tools sent inputs in English, the API returned a 400 Bad Request error. A similar behavior was observed in another healthcare application, the Cancer Registry of Norway, which also processes cancer case data in Norwegian isaku2023cost .
@@ 6.2 Automation Level
Each REST API testing tool supports different levels of automation. Our observations are based on running these tools locally using their provided Java or Python executables, rather than containerized setups like Docker, which may result in different user experiences. A common input for all tools is an OpenAPI specification in Swagger or JSON format, which can be easily converted between formats; for instance, RestTestGen provides automated support for this conversion. We found EvoMaster and Schemathesis straightforward to use in black-box mode, requiring only the OpenAPI specification and basic parameter settings. In contrast, tools like RestTestGen, RESTler, and RESTest require additional configurations. RestTestGen requires defining configuration files for each API and testing technique, including tool-specific and API-specific settings. RESTler necessitates the installation of Microsoft .NET and running a Compile phase for each API individually, which initially resulted in errors that required manual debugging to resolve configuration issues. Similarly, RESTest requires manual setup of API properties and test configuration files; although it can generate a configuration skeleton, it must be tailored manually. While these tools necessitate manual effort for initial configuration and setup, the subsequent testing process is automated. However, in a DevOps context, the manual effort required for API-related configurations can increase considerably due to the continuous evolution of APIs. A potential solution to alleviate manual effort involves utilizing automated tools that can infer REST API specifications as their implementation undergoes evolution. For example, a tool like RESTSpecIT decrop2024you can be used for this purpose.
@@ 6.3 Regression Test Selection, Minimization, and Prioritization
Numerous techniques are presented in the literature for prioritization, minimization, and selection of the test suite in the context of regression testing yoo2012regression ; greca2023state . However, to the best of our knowledge, their applicability to prioritize, minimize, and select regression test cases for REST APIs remains undetermined. Proposing new techniques for regression test suite prioritization, minimization, or selection falls outside the scope of this paper. Therefore, in our experiments, we used retest-all strategy, which involves running all the tests generated for the initial release on all subsequent releases. Given the daily or weekly release routine within the DevOps process, executing all test cases becomes impractical due to the significant overhead of API calls. For IoT applications’ REST APIs dependent on various third-party applications, the frequent execution of numerous tests is infeasible due to the API call limit imposed by these external applications. Furthermore, in DevOps environments, API specifications typically evolve along with the implementation. As a result, some test cases become obsolete, which requires strategies such as test case minimization to systematically execute tests on subsequent releases. A potential area for future research could focus on devising strategies for prioritization, minimization, or selection of test cases specifically for the regression testing of REST APIs of IoT-based applications. The regression testing outcomes presented in this work could serve as a foundation for further research. These findings could also be used to refine or customize existing methods for test selection, minimization, and prioritization within the REST API testing context.
@@ 6.4 REST API Regression Testing Tool
6.5 Continuous Regression Testing in DevOps
Typically, unintended faults or behavioral changes identified through regression testing are investigated and fixed before moving to the next release. However, our experiment was not designed in this way, as it would have required continuous involvement from industry practitioners to debug and fix issues for each release. This level of involvement was not feasible within the scope of our project due to the significant time and resources required, including specialized personnel (e.g., software engineers and testers), additional medical devices, and dedicated computing infrastructure. In our study, we shared all experimental results with the Oslo City team and their industry partner upon completion of the experiment and analysis. Due to organizational policies and confidentiality rules, we received general positive feedback indicating that the results were insightful, without confirmation of whether specific regressions or faults were present or fixed. Conducting a study with continuous regression testing in a DevOps context would require active practitioners’ engagement throughout the experiment. It would also require applying techniques such as regression test selection, minimization, or prioritization to systematically run tests. Therefore, such a study represents a promising direction for future research.
6.6 Findings Comparison with Existing Studies
As our study focuses on black-box REST API testing, we compare our experimental findings and conclusions with those reported in previous black-box testing studies. The work by Corradini et al. corradini2021empirical presented a comparison of four tools (RestTestGen, RESTler, bBOXRT, and RESTest) based on robustness and coverage. They found RESTler to be the most robust and RESTest the least robust. Although our analysis did not focus on robustness, we encountered no execution failures beyond initial configurations. Regarding coverage, their study reported that RestTestGen achieved high coverage, whereas our results showed comparable coverage across RestTestGen, RESTler, and RESTest, with RESTest slightly lower. Their general conclusions were that (i) RESTler is a mature but time-consuming tool, (ii) RestTestGen is better suited for time-constrained testing, (iii) combining multiple input generation techniques could be useful, and (iv) achieving high input coverage is easier than output coverage. Our study only supports their conclusion regarding high input coverage. However, we did not observe other conclusions, as we did not assess the time consumption of each tool or explore combinations of input generation techniques.
In the work by Kim et al. kim2022automated , 10 tools were compared in black-box and white-box modes based on code coverage, error responses, and practical implications. Although their findings are based on code coverage, ours focuses on API coverage. A common finding with our work is that erroneous requests are sometimes accepted, which, in our case, enabled identifying faults. Their conclusions highlighted that (i) sample input parameter values can improve test generation, (ii) natural language processing (NLP) can be used to interpret parameter descriptions, (iii) black-box tools could not achieve high coverage, and (iv) Schemathesis-PT and EvoMaster-BB outperformed other tools. Our study does not report the first two conclusions. For the latter two, our findings differ: all black-box tools achieved over 82% coverage, and the effectiveness of each tool varied by context, such as RESTest outperformed in fault detection and EvoMaster-BB outperformed in regression identification. A possible reason for these differences is that we used more recent versions of the tools than those evaluated in kim2022automated , and the newer or upgraded features may have contributed to improved coverage and fault detection.
In the work by Zhang et al. zhang2023open , seven tools were evaluated in black-box and white-box modes based on line coverage, time budget, faults, coverage-fault correlation for white-box testing, and current challenges. Their black-box findings indicated that EvoMaster-BB and Schemathesis-PT achieved the highest line coverage, longer time budgets yielded better results, and each tool detected only a few fault types. Our study did not analyze line coverage or various time budgets, but found RESTest tests effective in identifying most types of faults. Their overall conclusions were that (i) many tools lack robustness and crash frequently, (ii) API schemas may contain faults, (iii) a high number of interactions with databases, (iv) the need for mocking external live services, and (v) NLP can be used to infer constraints on string input parameters. In comparison, none of these conclusions are reported in our study, as our analysis focused specifically on black-box testing. For instance, we did not study aspects such as robustness, API specification faults, or mocking of external services. Moreover, some of their conclusions were derived from white-box testing contexts, such as database interactions.
@@ 7 Conclusion
The results revealed that EvoMaster-BB generated the highest number of test cases compared to all other tools. On the other hand, RestTestGen-MAST did not generate any tests, indicating that none of the APIs have security vulnerabilities. In terms of API failures, all tools generated a higher number of client-side errors than server-side errors. Among all, EvoMaster-BB generated tests led to the highest number of API failures. The tests produced by all tools led to the identification of 18 unique potential faults across 14 releases, with RESTest-CBT and RESTest-RT being the primary contributors to these faults. As for API test coverage, the results indicated that all tools attained a coverage level of 82-84%, with the lowest coverage observed in status code and status code class coverages. The regression results showed that tests generated by EvoMaster-BB led to finding the highest number of regressions. In total, 23 unique potential regressions were identified in all 14 releases. Finally, the cost analysis demonstrated that each tool generates a significant number of tests, often leading to numerous duplicate failures, thereby resulting in a cost overhead exceeding 70%.